# Ringdown waves from hairy black holes

### 16 Mar 2026 v1

@@ 6 Summary and discussions
In this paper, we derived the expressions of the QNM frequencies for stationary hairy black holes by using the correspondence between the circular null geodesics and QNMs. We modeled the deviation from the corresponding vacuum solution for the Einstein equations as perturbations in the form of an anisotropic fluid. Then, we utilized the relation between the eikonal limit of QNMs and null circular orbits. We explicitly obtained QNM formulas expressed in terms of the state parameters of the hair field. Our formulas provide a systematic approach to the analysis of QNMs for hairy black holes without assuming any specific underlying theory or model. We also examined the relationship between the state parameters and the QNM frequencies, including possible violations of the energy conditions. For the first two static models discussed in this work, the small-deviation approximation from the vacuum solution is incompatible with the dominant energy condition (DEC). This implies that, if the corresponding deviation is confirmed through the QNM observations, we need to consider a matter field violating the DEC or any other modification in the gravitational wave emission mechanisms. Since the violation of the DEC implies a superluminal flux, it might be healthier to consider a modified gravity theory than to rely on an exotic matter field. WEC and SEC are satisfied for all our examples, therefore solutions still maintain a positive energy density, and gravity is still locally attractive in both timelike and null geodesics.
In our model, we considered the rotating Bardeen, Hayward, and Kiselev metric constructed by applying Janis-Newman method to their static counterpart. Then we restricted our attention to the UCOP on the equatorial plane and obtain angular frequency and Lyapunov exponent to get the QNM information. The correspondence between the QNM and the UCOP for the rotating case is not yet fully established, but for the UCOP on the equatorial plane its considered to be related to the $l=|m|$ mode of QNM [2].
@@
In the construction of our hairy black hole models, we considered the effects of matter fields as perturbative deviations from the Schwarzschild and the Kerr solutions. In particular, we assumed the simplest form of the equations of state (2.15) and focused only on the first order perturbative effects of the matter field. It should be noted that our formulas given in Sec.  2.1 (also the formula (2.14)) for the static case, and those in Sec. 4.1 (see e.g., (4.14), (4.15)) for the rotating case can apply for the general cases. It would therefore be interesting to extend our present analyses to higher order perturbative analyses, just like higher order WKB analyses, and to include more general equation of state, such as the polytrope type. Possible modifications of QNMs induced by matter fields or deviations from general relativity were also discussed by Cardoso et al. [23] and McManus et al. [24], who considered perturbative deformations of the effective potential in the wave equations. It would be interesting to compare their results with those obtained in our analysis.

### 18 Mar 2026 v2

# Report number: NU-QG-15 Ringdown waves from hairy black holes

@@ 6 Summary and discussions
In this work, we developed a unified perturbative framework to study QNM frequencies of stationary hairy black holes by exploiting the correspondence between the unstable circular null geodesics and QNMs. We modelled deviations from the corresponding vacuum solutions of the Einstein equations as perturbations in the form of an anisotropic fluid. Using the relation between the eikonal limit of QNMs and unstable null circular orbits, we derived explicit formulas for the QNM frequencies in terms of the state parameters of the hair field. Our formulas provide a systematic approach to the analysis of QNMs for hairy black holes without assuming any specific underlying theory or model. We also examined the relationship between the state parameters and the QNM frequencies, including possible violations of the energy conditions.
In the static case, we showed that the QNM shifts can be expressed directly in terms of the matter distribution and its equation of state, with the Lyapunov exponent receiving an additional explicit contribution from the tangential pressure. In particular, for the Bardeen and Hayward black hole models, the small-deviation approximation from the vacuum solution is incompatible with the dominant energy condition (DEC). This implies that, if the corresponding deviation is confirmed through the QNM observations, we need to consider a matter field violating the DEC or any other modification in the gravitational wave emission mechanisms. Since the violation of the DEC implies a superluminal flux, it might be more physically reasonable to consider a modified gravity theory than to rely on exotic matter fields. By contrast, for the Kiselev case, the compatibility with the energy conditions depends on the parameters $k$ and $w_{q}$, and all of the standard energy conditions can be satisfied within suitable parameter ranges even in the small-$|k|$ regime. WEC and SEC are satisfied for all our examples, solutions still maintain a positive energy density, and gravity is still locally attractive in both timelike and null geodesics.
In the rotating case, the QNM components are separated into co-rotating and counter-rotating branches according to the equatorial UCOP structure. This shows that rotation does not merely shift the static result, but introduces an asymmetry in how the hair modifies the oscillation frequency and damping rate, with both branches reducing smoothly to the static case in the non-rotating limit. We considered the Bardeen, Hayward, and Kiselev metrics obtained by applying the Janis–Newman method to their static counterparts. We then restricted our attention to UCOPs on the equatorial plane and derived the corresponding angular frequencies and Lyapunov exponents, from which the QNM information can be inferred. Moreover, because the effective energy density and principal pressures in the co-rotating frame are the same as in the corresponding static case, the energy-condition analysis performed for the static models remains applicable to the rotating case as well. Although the correspondence between QNMs and UCOPs is not fully established in the rotating case, equatorial UCOPs are expected to be associated with the $\ell=|m|$ mode of QNMs [4].
@@
In the construction of our hairy black hole models, we considered the effects of matter fields as perturbative deviations from the Schwarzschild and the Kerr solutions. In particular, we assumed the simplest form of the equations of state (2.15) and focused only on the first-order perturbative effects of the matter field. It should be noted that our formulas given in Sec. 2.1 (also the formula (2.14)) for the static case, and those in Sec. 4.1 (see e.g., (4.14), (4.15)) for the rotating case can apply for the general cases. It would therefore be interesting to extend our present analyses to higher-order perturbative analyses, just like higher-order WKB analyses, and to include more general equations of state, such as the polytrope type. Possible modifications of QNMs induced by matter fields or deviations from general relativity were also discussed by Cardoso et al. [5] and McManus et al. [20], who considered perturbative deformations of the effective potential in the wave equations. It would be interesting to compare their results with those obtained in our analysis.
Overall, our results suggest that black hole ringdown signals can probe not only the spacetime geometry, but also the physical nature of the matter environment and possible deviations from vacuum general relativity in the vicinity of black holes.